# Temporal Sparse Autoencoders: Leveraging the Sequential Nature of Language for Interpretability

### v1

@@ 4 Experimental Evaluation
In this section, we rigorously evaluate Temporal SAEs. In Sec. 4.2, we evaluate our T-SAE’s recovery and disentanglement of semantic, contextual, and syntactic content, in Sec. 4.3, we report results on standard SAE evaluation metrics, in Sec. 4.4, we measure various aspects of temporal consistency, in Sec. 4.5, we present a case study of how SAEs can help understand alignment data and steer models, and finally in Sec. 4.6, we provide results on various ablation studies.
@@ 4.1 Implementation Details
Models and Datasets. We conduct all experiments on Pythia-160m [31] and Gemma2-2b [32]. We compare against baselines of BatchTopK SAEs [33], Matryoshka SAEs [11], and the model latents themselves when applicable. All models are trained and tested on subsets of the Pile [34]. All probing evaluations are done on MMLU [35], Wikipedia [36], and FineFineWeb [37].
Hyperparameters. We train Pythia SAEs on layer 8 and Gemma SAEs on layer 12. All SAEs are trained with the BatchTopK activation (batch-$k$-sparsity of 20), a dimensionality of 16k features, and the auxiliary loss from [11, 38]. These layers and hyperparameters are chosen to allow for comparability with pretrained and evaluated SAEs on Neuronpedia [39]. Temporal and Matryoshka SAEs are trained with 20%-80% feature splits, where for T-SAEs the 20% are the high-level features. We use a regularization parameter of 1.0 on the temporal loss for all T-SAEs. For ablations on hyperparameters, please see Section 4.6.
@@ 4.2 Probing for Semantics, Context, and Syntax
Figure 3: Accuracy of probes trained on SAE decompositions for Gemma2-2b, as well as probes trained directly on model latents (orange), with semantic labels (right), contextual labels (middle), and syntactic labels (right) with varying levels of probe sparsity (setup from [40]). T-SAEs significantly outperform baseline SAEs for semantics and context.
@@
To understand the types of features that the temporal loss encourages SAEs to learn, we evaluate the ability of T-SAEs to recover different types of linguistic information, namely semantic, contextual, and syntactic. We provide qualitative visualizations of these results in Figure 2 and quantitative probing results in Figure 3.
@@
In Figure 2, we present TSNE [41] visualizations of the Pythia-160m SAE activations for various questions taken from the MMLU dataset, which we color by the semantic content of the question, the context of which sequence a token came from, and the syntactic information for each token. In particular, we encode 20 tokens from each question into the SAE’s feature space, and use TSNE as a dimensionality reduction and visualization method to understand how the SAE embeddings are clustered. We use the question category, such as “High School European History" or “Professional Medicine" as proxies for the semantic information. The contextual information simply refers to the question ID of each token, meaning tokens from the same question or context should have the same color. We use an open-source NLP library, spaCy [42], to retrieve part-of-speech labels for each token as a proxy for their syntactic content.
@@
We find that the activations of high-level features from T-SAEs cluster strongly according to semantic content (top left) and contextual content (top middle), meaning tokens from the same question are represented similarly, as well as questions from the same topic. On the other hand, activations of the low-level features seem to be more syntactic, with stronger clusters for parts-of-speech (middle right). Matryoshka SAE embeddings prioritize syntactic information strongly over semantic and contextual information, with clear clustering for part-of-speech (bottom right) and minimal clustering for the other two labels.
Probing Validation.
@@
In order to validate these visual results, we probe the SAE activations for these same labels, using both k-sparse probing from Kantamneni et al. [40] as well as normal Logistic Regression probes on the full activations. We train probes on k=1,5,10,20k={1,5,10,20} features, where we select the features by comparing the mean activations of the positive and negative examples for each class from the train set. Note that dense probes trained on SAE activations have a dimensionality of 16k, whereas dense probes trained directly on the model’s residual stream have a dimensionality of 768 for Pythia-160m and 2304 for Gemma2-2b. We find that these results quantitatively reflect the qualitative results from the TSNE plots, with T-SAEs outperforming the baseline SAEs significantly for semantic and contextual labels, with little-to-no performance drop for syntactic information. Probing results for two more datasets, Wikipedia and FineFineWeb, as well as for Pythia-160m, are in Appendix A.2, with the same trends across all models and datasets.
Feature Disentanglement.
@@
While we see that Temporal SAEs recover more semantic and contextual information, can this behavior be attributed to the high-level features alone? Indeed, we observe specialization between high- and low-level features in Figure 2, where the high-level features exhibit semantic and contextual structure, whereas low-level features exhibit syntactic structure. To further characterize this behavior, we report probing accuracy on each of the feature splits separately (see Appendix A.1), which confirms this same specialization. Interestingly enough, despite not performing the reconstruction task on their own due to the Matryoshka training loss, the low-level features are able to recover syntactic information. In contrast, for Matryoshka SAEs, across all tasks, performance can be attributed almost entirely to the high-level feature split, with the low-level features being significantly less predictive for semantics, context, and syntax, indicating a lack of specialization of high- and low-level features in the baseline Matryoshka SAEs.
@@ 4.3 SAE Evaluation
In Table 1, we provide results for various standard evaluations of SAEs to ensure that temporal consistency does not significantly degrade reconstruction quality. We define our metrics as such. Fraction Variance Explained (FVE): The fraction of the SAE input data’s total variance that is successfully captured by the SAE decomposition, or $1-\mathrm{Var}(x-\hat{x})/\mathrm{Var}(x)$. Cosine Similarity (Cos Sim): The cosine similarity between the SAE inputs and outputs. Fraction Alive: The proportion of SAE features that activate at least once on the test data. Activation Smoothness: For each sequence $s$, we filter for active features (features that fire at least once over the sequence). We compute $\Delta_{s}=\frac{1}{n^{\prime}}\sum_{i=1}^{n^{\prime}}\max_{t\in{[1...T]}}|\mathbf{f}_{i}(\mathbf{x}_{t})-\mathbf{f}_{i}(\mathbf{x}_{t-1})|/\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\|_{2}$ the average max absolute change over the $n^{\prime}$ active features normalized by the change in model latents [^3] This can be viewed as the average per-feature Lipschitz constant across sequences.. Finally, we average over multiple sequences to get a smoothness score, $S=\frac{1}{n}\sum_{s}^{n}\Delta_{s}$.
Automated Interpretability (Autointerp) Score: Score for how correct feature explanations are. We use SAEBench [43] to generate and score feature explanations with Llama3.3-70B-Instruct. For each latent, the LLM generates potential feature explanations based on a range of activating examples. Then, we collect activating and non-activating examples and ask a judge (also Llama3.3-70B-Instruct) to use the feature explanation to categorize examples and score its performance. [^4] Note that both the generation and scoring phase are highly noisy and dependent on the LLM judge.
@@
Table 1: Core Performance Metrics. We report smoothness on the high-level split (H) and standard deviations for autointerpretability scores.
FVE ($\uparrow$) Cos Sim ($\uparrow$)
Fraction
Alive ($\uparrow$)
Activation
Smoothness ($\downarrow$)
Autointerp
Score ($\uparrow$)
Pythia 160m
Temporal SAE 0.94 0.93 0.87 0.09 (H) 0.81 $\pm$ 0.17
Matryoshka SAE 0.95 0.94 0.89 0.12 0.83 $\pm$ 0.16
BatchTopKSAE 0.95 0.94 0.84 0.13 0.85 $\pm$ 0.15
Gemma 2-2b
Temporal SAE 0.75 0.88 0.78 0.10 (H) 0.83 $\pm$ 0.15
Matryoshka SAE 0.75 0.89 0.76 0.14 0.83 $\pm$ 0.16
BatchTopKSAE 0.76 0.89 0.66 0.13 0.83 $\pm$ 0.16
Figure 4: Top 8 most active Gemma2-2b Temporal SAE features over a concatenated sequence of text. T-SAE features exhibit clear phase transitions between sequences, are relatively smooth, and have explanations relevant to the semantic content of each component sequence.
@@ 4.4 Temporal Consistency
Table 1 indicates that T-SAE features are smoother and more consistent than baselines. To explore this further, we visualize the top feature activations of T-SAEs over long sequences of text (Figs. 1, 4). In Figure 4, we concatenate four sequences of text: a biology question (MMLU), a letter from Charles Darwin (Project Gutenberg), an article on Animal Farm (Wikipedia), and a mathematics question (MMLU), and interpret them with a T-SAE. We take the top-8 most active features across the sequence and plot their activations for each token. The T-SAE features have clear phase transitions between texts, with different features activating and deactivating for specific sequences. Furthermore, within a sequence, these features are relatively smooth, without spiky, high-frequency changes in activation from token to token. We can even detect periodic behavior in the biology question, corresponding to each multiple-choice answer option. We note that there is an interesting leakage of features that continue to fire in later, semantically unrelated sections of the spliced sequence, highlighting that the model may retain past context and that the T-SAEs are able to identify this information rollover. Figure 1 conducts a similar analysis over Newton’s Principia (Project Gutenberg), a genetics question (MMLU), and the Bhagavat Gita (Project Gutenberg) and compares the top features provided by the Temporal SAE to those generated by a baseline Matryoshka SAE. We find similar consistency and smoothness over sequences with clear transitions between sequences for the T-SAE, whereas the top Matryoshka SAE features activate across all three sequences without differentiating them, and are much noisier and high-frequency, fluctuating across tokens. Finally, we interpret these features using automated interpretability and report their explanations in the figure as annotations about the activations. We find that the labels reflect the true underlying semantic content present in each component sequence, with the ‘Animal Farm’ Wikipedia article activating a feature for “Historical literature and academic writing" in Figure 4 and the Bhagavat Gita excerpt in Figure 1 activating a “Worship and spiritual practices" feature.
@@
Sequence-level Interpretability. Figure 1 and prior work [15] demonstrate how baseline SAE features are “dense,” in that they fluctuate frequently over a sequence. In doing so, they can only provide human-interpretable explanations at the token level; as soon as you zoom out, you get an overwhelming set of activating features with very little structure or parseability. The smoothness of T-SAEs unlock a sequence-level understanding of data which was previously much harder, if not impossible, to parse from baseline SAE explanations (Figs. 1, 4).
Figure 5: Left. We study the features describing the HH-RLHF [44] dataset given by our T-SAE and a Matryoshka SAE provided by Neuronpedia. The Matryoshka SAE appears to find more random features, whereas the T-SAe is able to capture safety-relevant features. Additionally, the T-SAE sheds light on a potential length spurious correlation in the dataset, where rejected completions are longer than chosen completions. Upon investigating the dataset, we find this is true with statistical significance, highlighting the potential of the high-level features for data filtering, steering, and targeted finetuning interventions. Right. We find that high-level features Pareto dominate baselines in their ability to steer LLMs. Specifically, they are better at preserving coherence while successfully changing the semantics of model generation.

### v2

@@ 4 Experimental Evaluation
In this section, we evaluate Temporal SAEs. In Sec. 4.2, we evaluate our Temporal SAE’s recovery and disentanglement of semantic, contextual, and syntactic content, in Sec. 4.3, we report results on standard SAE evaluation metrics, in Sec. 4.4, we measure various aspects of temporal consistency, in Sec. 4.5, we present a case study of how SAEs can help to uncover safety-relevant concepts and mechanisms in LLMs, and finally in Sec. 4.6, we provide results on various ablation studies.
@@ 4.1 Implementation Details
Models and Datasets. We conduct all experiments on Pythia-160m (Biderman et al., 2023) and Gemma2-2b (Team et al., 2024). We compare against baselines of BatchTopK SAEs (Bussmann et al., 2024), Matryoshka SAEs (Bussmann et al., 2025), and the model latents themselves when applicable. All models are trained and tested on the Pile (Gao et al., 2020). All probing evaluations are done on MMLU (Hendrycks et al., 2020), Wikipedia Wikipedia (2004), and FineFineWeb (M-A-P et al., 2024).
Hyperparameters. We train Pythia SAEs on layer 8 and Gemma SAEs on layer 12. All SAEs are trained with the BatchTopK activation (kk=20), 16k features, and the auxiliary loss from (Bussmann et al., 2025; Gao et al., 2024). These models, layers, and hyperparameters are chosen to allow for comparability with pretrained and evaluated SAEs on Neuronpedia (Lin, 2023). Temporal and Matryoshka SAEs are trained with 20%-80% feature splits, where for Temporal SAEs the 20% are the high-level features. We use a regularization parameter of 1.0 on the temporal loss for all Temporal SAEs. For ablations on hyperparameters, see Section 4.6.
@@ 4.2 Probing for Semantics, Context, and Syntax
Figure 3: Accuracy of probes trained on SAE decompositions for various SAEs trained on Gemma2-2b, as well as probes trained directly on model latents (orange), with semantic labels (right), contextual labels (middle), and syntactic labels (right) with varying levels of probe sparsity (setup from (Kantamneni et al., 2025)). Dense probes are trained on all features.
@@
To understand the types of features that the temporal loss encourages SAEs to learn, we evaluate the ability of Temporal SAEs to recover different types of linguistic information, namely semantic, contextual, and syntactic. We provide qualitative visualizations of these results in Figure 2 and quantitative probing results in Figure 3.
@@
In Figure 2, we present TSNE visualizations of the Pythia-160m SAE activations for various questions taken from the MMLU dataset, which we color by the semantic content of the question, the context of which sequence a token came, and the syntactic information for each token. In particular, we encode the last 20 tokens from each question into the SAE’s feature space, and use TSNE as a dimensionality reduction and visualization method to understand how the SAE embeddings are clustered. We use the question category, such as “High School European History” or “Professional Medicine” as proxies for the semantic information. The contextual information simply refers to the question ID of each token, meaning tokens from the same context should have the same color. We use an open-source NLP library, spaCy (Honnibal et al., 2020), to retrieve part-of-speech labels for each token as a proxy for their syntactic content.
@@
We find that the activations of high-level features from Temporal SAEs cluster strongly according to semantic content (top left) and contextual content (top middle), meaning tokens from the same question are represented similarly, as well as questions from the same topic. On the other hand, activations of the low-level features seem to be more syntactic, with stronger clusters for parts-of-speech (middle right). Matryoshka SAE embeddings prioritize syntactic information strongly over semantic and contextual information, with clear clustering for part-of-speech (bottom right) and minimal clustering for the other two labels.
@@
Probing Validation. In order to validate these visual results, we probe the SAE activations for these same labels, using both k-sparse probing from Kantamneni et al. (2025) as well as normal Logistic Regression probes on the full activations. We train probes on k=1,5,10,20k={1,5,10,20} features, where we select the features by comparing the mean activations of the positive and negative examples for each class from the train set. Note that dense probes trained on SAE activations have a dimensionality of 16k whereas dense probes trained directly on the model’s residual stream have a dimensionality of 768 for Pythia-160m and 2304 for Gemma2-2b. We find that these results quantitatively reflect the qualitative results from the TSNE plots, with Temporal SAEs outperforming the baseline SAEs significantly for semantic and contextual labels, with little-to-no performance drop for syntactic information. Probing results for two more datasets, Wikipedia and FineFineWeb, as well as for Pythia-160m, are in Appendix C.5, with the same trends across all models and datasets.
@@
Feature Disentanglement. While we see that Temporal SAEs recover more semantic and contextual information, can this behavior be attributed to the high-level features alone? Indeed, we observe specialization between high- and low-level features in Figure 2, where the high-level features exhibit semantic and contextual structure, whereas low-level features exhibit syntactic structure. Similarly, our smoothness metrics (Table 1) show that T-SAE high-level features are smoother than baselines and T-SAE low-level features are less smooth, likewise providing evidence of separation. To further validate this behavior, we measure probing accuracy on each feature split separately and find high- and low-level task specialization (Appendix C.1). Interestingly, we find that the low-level features are able to recover syntactic information on their own, despite only be trained to reconstruct the residual left by the high-level split. In contrast, for Matryoshka SAEs, across all tasks, performance can be attributed almost entirely to the high-level feature split, with the low-level features being significantly less predictive for semantics, context, and syntax, indicating a lack of disentanglement in baselines.
Disentanglement allows users to (1) attribute behavior to specific types of features and (2) more precisely control the types of features they care about. For example, in domains such as safety monitoring, users may care more about the semantics of the output (e.g. whether the output contains sexually-explicit content), whereas in poetry or coding tasks, users may care about syntax and local features (e.g. vowel sounds or closing brackets).
@@ 4.3 SAE Evaluation
In Table 1, we provide results for various standard evaluations of SAEs to ensure that temporal consistency does not significantly degrade reconstruction quality. Fraction Variance Explained (FVE): The fraction of the SAE input data’s total variance that is successfully captured by the SAE decomposition, or $1-\mathrm{Var}(x-\hat{x})/\mathrm{Var}(x)$. Cosine Similarity (Cos Sim): The cosine similarity between the SAE inputs and outputs. Fraction Alive: The proportion of SAE features that activate at least once on the test data. Smoothness: For each sequence $s$, we filter for active features (features that fire at least once over the sequence). We compute $\Delta_{s}=\frac{1}{n^{\prime}}\sum_{i=1}^{n^{\prime}}\max_{t\in{[1...T]}}|\mathbf{f}_{i}(\mathbf{x}_{t})-\mathbf{f}_{i}(\mathbf{x}_{t-1})|/\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\|_{2}$ the average max absolute change over the $n^{\prime}$ active features normalized by the change in model latents [^3] This can be viewed as the average per-feature Lipschitz constant across sequences.. Finally, we average over multiple sequences to get a smoothness score, $S=\frac{1}{n}\sum_{s}^{n}\Delta_{s}$. Automated Interpretability (Autointerp) Score: Score for how correct feature explanations are. We use SAEBench (Karvonen et al., 2025) to generate and score feature explanations with Llama3.3-70B-Instruct. For each latent, the LLM generates potential feature explanations based on a range of activating examples. Then, we collect activating and non-activating examples and ask a judge (also Llama3.3-70B-Instruct) to use the feature explanation to categorize examples and score its performance. [^4] Note that both the generation and scoring phase are highly noisy and dependent on the LLM judge.
@@
Table 1: Core Performance Metrics. We report smoothness on feature splits when applicable and standard deviations for autointerpretability scores.
FVE ($\uparrow$) Cosine Sim ($\uparrow$) Fraction Alive ($\uparrow$) Smoothness Autointerp Score ($\uparrow$)
Full High Low
Pythia 160m Temporal SAE 0.94 0.93 0.87 0.12 0.09 0.17 0.81 $\pm$ 0.17
Matryoshka SAE 0.95 0.94 0.89 0.12 0.12 0.13 0.83 $\pm$ 0.16
BatchTopKSAE 0.95 0.94 0.84 0.13 - - 0.85 $\pm$ 0.15
Gemma 2-2b Temporal SAE 0.75 0.88 0.78 0.13 0.10 0.15 0.83 $\pm$ 0.15
Matryoshka SAE 0.75 0.89 0.76 0.14 0.15 0.12 0.83 $\pm$ 0.16
BatchTopKSAE 0.76 0.89 0.66 0.13 - - 0.83 $\pm$ 0.16
@@ 4.4 Temporal Consistency
Table 1 indicates that Temporal SAE features are smoother and more consistent than baselines. To explore this further, we visualize the top feature activations of Temporal SAEs over long sequences of text (Figs. 1, 4). In Figure 4, we concatenate four sequences of text: a biology question (MMLU), a letter from Charles Darwin (Project Gutenberg), an article on Animal Farm (Wikipedia), and a mathematics question (MMLU), and interpret them with a Temporal SAE. We take the top-8 most active features across the sequence and plot their activations for each token. The Temporal SAE features have clear phase transitions between texts, with different features activating and deactivating for specific sequences. Furthermore, within a sequence, these features are relatively smooth, without spiky, high-frequency changes in activation from token to token. We can even detect periodic behavior in the biology question, corresponding to each multiple-choice answer option. We note that there is an interesting leakage of features that continue to fire in later, semantically unrelated sections of the spliced sequence, highlighting that the model may retain past context and that the Temporal SAEs are able to identify this information rollover. Figure 1 conducts a similar analysis over Newton’s Principia (Project Gutenberg), a genetics question (MMLU), and the Bhagavat Gita (Project Gutenberg) and compares the top features provided by the Temporal SAE to those generated by a baseline Matryoshka SAE. We find similar consistency and smoothness over sequences with clear transitions between sequences for the Temporal SAE, whereas the top Matryoshka SAE features activate across all three sequences without differentiating them, and are much noisier and high-frequency, fluctuating across tokens. Finally, we interpret these features using automated interpretability and report their explanations in the figure as annotations about the activations. We find that the labels reflect the true underlying semantic content present in each component sequence, with the ‘Animal Farm’ Wikipedia article activating a feature for “Historical literature and academic writing” in Figure 4 and the Bhagavat Gita excerpt in Figure 1 activating a “Worship and spiritual practices” feature.
@@
Sequence-level Interpretability. Figure 1 and prior work (Sun et al., 2025) demonstrate how baseline SAE features are “dense,” in that many features fluctuate frequently over a sequence. In doing so, they can only provide human-interpretable explanations at the token level; as soon as you zoom out, you get an overwhelming set of activating features with very little structure or parseability. The smoothness of Temporal SAEs unlock a sequence-level understanding of data which was previously much harder, if not impossible, to parse from baseline SAE explanations (Figs. 1, 4).
Figure 4: Top 8 most active Gemma2-2b Temporal SAE features over a concatenated sequence of text. Temporal SAE features exhibit clear phase transitions between sequences, are relatively smooth, and have explanations relevant to the semantic content of each component sequence.